# Efficient Generation of Smooth Paths with Curvature Guarantees by Mollification

### v1

@@ IV-B Quasiconvexity under mollification
Definition 3 (Quasiconvex function)
@@
That is, $\gamma x+(1-\gamma)y\in S_{\alpha}$. Therefore $S_{\alpha}$ is a convex set, proving $f\circ g$ is quasiconvex. ∎
@@
Proposition 6 (Monotonicity and mollification)
@@
since $\varphi_{\varepsilon}$ is positive. The proof is identical for monotone decreasing functions. ∎
Figure 4: Mollification of stair-step function created using shifted Heaviside step functions. The blue solid line represents the (discontinuous) stair-step function $f$, while the red solid line represents its mollification using the mollifier of equation (1) and $\varepsilon=0.5$. Due to proposition 6, $F_{\varepsilon}$ cannot present overshoots or oscillations.
@@
Theorem 5
Suppose $f:\mathbb{R}\to\mathbb{R}$ is a quasiconvex function and let $\varphi$ be a mollifier satisfying 1; then, for all $\varepsilon>0$ the function $F_{\varepsilon}:=f*\varphi_{\varepsilon}$ is quasiconvex.
Proof:
@@
First, we claim that the set $V_{1}:=\{t\in(-\varepsilon,\varepsilon)\mid f(\gamma x+(1-\gamma)y-t)-f(x-t)>0\}$ is measurable and it has positive measure. By Lemma 4 we know that the composition of a quasiconvex function with an affine mapping is quasiconvex, hence measurable. Since the linear combination of two measurable function results in a measurable function, the set $V_{1}$ is measurable. Thus, if $\lambda(V_{1})$ were zero, then it would negligible in integration, which would imply that
@@
$\int_{(-\varepsilon,\varepsilon)\setminus V_{1}}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)\leq 0.$
@@
However, $f(y-t_{2})<f(x-t_{2})$ by equation (4), thus
@@
Using the same approach and considering equation (3) we have that
@@
but this leads to a contradiction. Therefore, the assumption that there exists an $\varepsilon>0$ for which $F_{\varepsilon}$ is not quasiconvex is false. That is, $F_{\varepsilon}$ is quasiconvex for any $\varepsilon>0$. ∎
Remark 4
@@
IV-C Closed paths
So far, we have worked with real-valued functions because we represent desired paths or trajectories as parametric functions, i.e., functions $f:\mathbb{R}\to\mathbb{R}^{n}$. Nevertheless, as previously noted, our results can be extended to functions $\phi:\mathbb{R}^{n}\to\mathbb{R}$, for which convexity-like properties can also be defined. This extension is of practical interest, since in some situations the desired path is a set of the form
$\mathcal{P}=\{x\in\mathbb{R}^{n}\mid\phi_{i}(x)=0,\quad i\in\{1,\dots,n-1\}\},$
for which previous results applying to each $\phi_{i}$. Nevertheless, there are situations in which the path arises from a parametric curve with compact domain, say $f:[a,b]\to\mathbb{R}^{n}$ where $\infty<a<b<\infty$, and it is preferable to work with $f$ as a whole rather than component by component. Since we lose some order-like properties in $\mathbb{R}^{n}$ relative to $\mathbb{R}$, we cannot directly extend convexity-like properties [^1] We refer to pure convexity properties.. Nonetheless, we still want to characterize (in advance) how mollification affects the complete function $f$. We now address the following question: given the original trajectory, does there exist a subset $U$ of $\mathbb{R}^{n}$ such that the mollified trajectory is contained in $U$ for any value of its parameter? The answer is affirmative, with $U$ being the convex hull of $f([a,b])$.
Definition 4
@@ IV-C Closed paths
$F_{\varepsilon}(t)=(f*\varphi_{\varepsilon})(t)=\left((f_{1}*\varphi_{\varepsilon})(t),\dots,(f_{n}*\varphi_{\varepsilon})(t)\right).$
@@
$\left\{F_{\varepsilon}(t)\mid t\in\operatorname{dom}f\right\}\subset\operatorname{co}\left\{f(t)\mid t\in\operatorname{dom}f\right\}.$
Proof:
Let $U=\operatorname{co}\{f(t)\mid t\in\operatorname{dom}f\}$. Define the extended real valued function $I_{U}:\mathbb{R}^{n}\to\mathbb{R}\cup\{-\infty,\infty\}$ as
@@
i.e., $I_{U}(F_{\varepsilon}(t))=0$ so $F_{\varepsilon}(t)\in U$. Since $t$ and $\varepsilon$ are arbitrary, the claim follows. ∎
Remark 5
As previously mentioned, it is common that the path is a continuous function defined in a compact subset of $\mathbb{R}$, i.e., $f:[a,b]\to\mathbb{R}^{n}$ with $\infty<a<b<\infty$. In such a case, we can extend the function $f$ to $\mathbb{R}$ as follows to get a new continuous function
@@
Note that when working with trajectories with compact domain, we must extend them as done in Remark 5.
Lemma 7
Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function and fix $\varepsilon>0$. Let $\bar{f}:[a-\varepsilon,b+\varepsilon]$ be its continuous extension as done in remark 5. The following two statements are true:
@@
Theorem 8
Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function and let $\varphi$ be a nonnegative mollifier. Fixed $\varepsilon>0$ let $\bar{f}:[a-\varepsilon,b+\varepsilon]\to\mathbb{R}^{n}$ be the continuous extension as in remark 5. Define $F:[a,b]\to\mathbb{R}^{n}$ as $F=\bar{f}*\varphi_{\varepsilon}$; then $L(F)\leq L(f)$.
Proof:
@@
thus, by definition of the supremum, $L(F)\leq L(f)$. ∎
Remark 6
Note that this does not imply that if $f$ is a geodesic between $f(a)$ and $f(b)$, then $F$ is also a geodesic between these points. This is because $F(a)\neq f(a)$ or $F(b)\neq f(b)$ may occur. While Theorem 6 and Remark 5 guarantee that $F([a,b])=(\bar{f}*\varphi_{\varepsilon})([a,b])\subset\operatorname{co}f([a,b])$, we cannot ensure that $F$ has the same starting and ending points as $f$. What Theorem 8 establishes is that by considering the actual starting and ending points of $F$, we can ensure that $L(F)\leq L(f)$. An example of this property is shown in figure 5. The original function, which is a linear interpolation of three points in $\mathbb{R}^{2}$ and whose domain is $[a,b]=[0,2]$, is extended to the domain $\varepsilon,2+\varepsilon]$ with $\varepsilon=0.5$ using the extension presented in remark 5. As it can be seen from each of its components, $F_{\varepsilon}(a)\neq f(a)$ and $F_{\varepsilon}(b)\neq f(b)$, and clearly $L(F_{\varepsilon})\leq L(f)$ as theorem 8 states. Finally note that, while $f_{2}$ can be considered a geodesic between the points $(0,0)$ and $(2,2)$ in $\mathbb{R}^{2}$ using the Euclidean norm, $(f_{2}*\varphi_{\varepsilon})$ is not a geodesic between those two points.
Figure 5: Visual representation of remark 6. The left picture represents the original path $f$ as a blue solid line, and as a red solid line its mollification $F_{\varepsilon}=(f_{1}*\varphi_{\varepsilon},f_{2}*\varphi_{\varepsilon})$ where $\varphi$ is defined in equation (1) and $\varepsilon=0.5$. The middle picture represents the same information but for the first component of the function and its mollification, while the right picture represents the same information but for the second component.
IV-D The effect of reparametrization and mollification
Suppose the desired path is encoded using a continuous function $f:[a,b]\to\mathbb{R}^{n}$ and for a given $\varepsilon>0$ we consider its mollification with parameter $\varepsilon$. As we have seen, we first need to extend the function $f$ to $\bar{f}:[a-\varepsilon,b+\varepsilon]$ as is done in Remark 5 or in Theorem 8. It may be of interest to reparametrize the curve so it is normalized, i.e., to find a function $g:[-\varepsilon,1+\varepsilon]\to[a-\varepsilon,b+\varepsilon]$, and consider $\bar{f}\circ g*\varphi_{\varepsilon}:[0,1]\to\mathbb{R}^{n}$, as the desired mollified path. How does the parameter of the mollification change under these conditions? That is, does there exist a $\eta=\eta(\varepsilon)>0$ such that
@@ IV-D The effect of reparametrization and mollification
V Curvature guarantees from a sequence of waypoints
@@
V-A The case of three points forming two segments
@@
Definition 6 (Two line segments function)
@@
For $r\in\mathbb{N}$ we define $\tilde{P}_{r}:=P_{r}-P_{r-1}$.
Note that both $f$ and $\bar{f}$ are continuous functions, hence locally integrable, and from Theorem 1 we have that if $F_{\varepsilon}:=\bar{f}*\varphi_{\varepsilon}$ where $\varphi$ is any mollifier and $\varepsilon>0$ then $F_{\varepsilon}^{\prime}=\bar{f}*\varphi_{\varepsilon}^{\prime}$. Fix $t\in\mathbb{R}$, and note that
$\int_{\mathbb{R}\setminus\{t-1\}}\bar{f}^{\prime}(t-s)\varphi_{\varepsilon}(s)\,\mathrm{d}s$ $=\int_{\mathbb{R}\setminus\{1\}}\bar{f}^{\prime}(s)\varphi_{\varepsilon}(s)\,\mathrm{d}s$
$=\int_{\mathbb{R}}\bar{f}^{\prime}(s)\varphi_{\varepsilon}(s)\,\mathrm{d}s,$
because $\lambda(\{1\})=0$, so we could take any function $g$ that is equal to $\bar{f}^{\prime}$ in $\mathbb{R}\setminus\{1\}$ but is defined in $t=1$, and we would reach that $\int_{\mathbb{R}}g(t-s)\varphi_{\varepsilon}(s)=\int_{\mathbb{R}\setminus\{t-1\}}\bar{f}^{\prime}(t-s)\varphi_{\varepsilon}(s)\,\mathrm{d}s$. That is, we can evaluate $(\bar{f}^{\prime}*\varphi_{\varepsilon})(1)$ even if $\bar{f}^{\prime}(1)$ is not defined. Exploiting the properties of the compact support of the mollifiers
$\int_{\mathbb{R}}\bar{f}^{\prime}(s)\varphi_{\varepsilon}(t-s)\,\mathrm{d}s$
$=\bar{f}(s)\varphi_{\varepsilon}(t-s)\mid_{s=t-\varepsilon}^{s=t+\varepsilon}-\int_{\mathbb{R}}\bar{f}(s)(-\varphi_{\varepsilon}^{\prime}(t-s))\,\mathrm{d}s$
$=\int_{\mathbb{R}}\bar{f}(s)\varphi_{\varepsilon}^{\prime}(t-s)\,\mathrm{d}s=(\bar{f}*\varphi_{\varepsilon}^{\prime})(t)=F_{\varepsilon}^{\prime}(t).$
This proves that in this case $(\bar{f}^{\prime}*\varphi_{\varepsilon})=(\bar{f}*\varphi_{\varepsilon}^{\prime})=F_{\varepsilon}^{\prime}$, which is a powerful result given that the derivative of $\bar{f}$ is constant almost everywhere.
Theorem 9
Let $f:[a,b]\to\mathbb{R}^{n}$ and consider any continuous extension of $f$ to $\mathbb{R}$, i.e., a function $\bar{f}:\mathbb{R}\to\mathbb{R}^{n}$ that is continuous, $\bar{f}\mid_{[a,b]}=f$ and it is differentiable almost everywhere in $\mathbb{R}$. If it happens that $f$ is differentiable almost everywhere in $[a,b]$ then it holds that for any $\varepsilon>0$ and mollifier
[…]
Proof:
The proof is immediate following the steps from the previous discussion. For a more general result involving distributions see [16, Sections 6.13 and 6.14]. ∎
Corollary 1
Let $f$ be the two lines segment function of equation (5) and $\bar{f}$ its extended version defined in equation (6). Then, for any $\varepsilon>0$ and any mollifier $\varphi$ it holds that
[…]
We can now exploit the results of Corollary 1 to obtain a formula for the curvature, as well as an upper bound.
V-A1 Computing the exact curvature
We know by the Corollary 1 that $F_{\varepsilon}^{\prime}=(\bar{f}*\varphi^{\prime}_{\varepsilon})=(\bar{f}^{\prime}*\varphi_{\varepsilon})$ everywhere. Consider a mollifier $\varphi$ and let $\varepsilon>0$. Given $t\in\mathbb{R}\setminus\{1\}$, note that
@@ V-A1 Computing the exact curvature
V-A2 Upper bounding the curvature
@@
Figure 6: Left plot represents of a three-point-two-segment function and its mollification. Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of equation (8). In this case the zone of maximum curvature corresponds to a point really close (or equal) to $t=1$. We have used the mollifier presented in Example 1.
Figure 7: Left plot represents of a three-point-two-segment function and its mollification. Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of equation (8). In this case the zone of maximum curvature does not correspond to the join of the two segments. We have used the mollifier presented in Example 1.
V-B The general case

### v2

@@ 4.2 Quasiconvexity under mollification
Definition 4.19 (Quasiconvex function).
@@
That is, $\gamma x+(1-\gamma)y\in S_{\alpha}$. Therefore $S_{\alpha}$ is a convex set, proving $f\circ g$ is quasiconvex.
@@
Proposition 4.22 (Monotonicity and mollification).
@@
since $\varphi_{\varepsilon}$ is positive. The proof is identical for monotone decreasing functions.
Figure 4: Mollification of stair-step function created using shifted Heaviside step functions. The blue solid line represents the (discontinuous) stair-step function $f$, while the red solid line represents its mollification using the mollifier (1) and $\varepsilon=0.5$. Due to Proposition 4.22, $F_{\varepsilon}$ cannot present overshoots or oscillations.
@@
Theorem 4.24.
Suppose $f:\mathbb{R}\to\mathbb{R}$ is a quasiconvex function and let $\varphi$ be a mollifier satisfying Assumption 1; then, for all $\varepsilon>0$ the function $F_{\varepsilon}:=f*\varphi_{\varepsilon}$ is quasiconvex.
Proof 4.25.
@@
First, we claim that the set $V_{1}:=\{t\in(-\varepsilon,\varepsilon)\mid f(\gamma x+(1-\gamma)y-t)-f(x-t)>0\}$ is measurable and it has positive measure. By Lemma 4.20 we know that the composition of a quasiconvex function with an affine mapping is quasiconvex, hence measurable. Since the linear combination of two measurable function results in a measurable function, the set $V_{1}$ is measurable. Thus, if $\lambda(V_{1})$ were zero, then it would be negligible in integration, which would imply that
@@
$\int_{(-\varepsilon,\varepsilon)\setminus V_{1}}\left[f(\gamma x+(1-\gamma)y-t)-f(x-t)\right]\varphi_{\varepsilon}(t)\mathrm{d}\lambda(t)\leq 0,$
@@
However, $f(y-t_{2})<f(x-t_{2})$ because of (4), thus
@@
Using the same approach and considering (3) we have that
@@
but this leads to a contradiction. Therefore, the assumption that there exists an $\varepsilon>0$ for which $F_{\varepsilon}$ is not quasiconvex is false. That is, $F_{\varepsilon}$ is quasiconvex for any $\varepsilon>0$.
Remark 4.26.
@@
4.3 Enclosure and length of paths
So far, we have worked with real-valued functions because we represent desired paths or trajectories as parametric functions, i.e., functions $f:\mathbb{R}\to\mathbb{R}^{n}$. Nonetheless, we still want to characterize (in advance) how mollification affects the complete function $f$, that is, treat it as a whole. We now address the following question: given the original trajectory, does there exist a subset $U$ of $\mathbb{R}^{n}$ such that the mollified trajectory is contained in $U$ for any value of its parameter? The answer is affirmative, with $U$ being the convex hull of $f(\mathbb{R})$.
Definition 4.27.
@@ 4.3 Enclosure and length of paths
$F_{\varepsilon}(t):=(f*\varphi_{\varepsilon})(t)=\left((f_{1}*\varphi_{\varepsilon})(t),\dots,(f_{n}*\varphi_{\varepsilon})(t)\right).$
@@
$\left\{F_{\varepsilon}(t)\mid t\in\mathbb{R}\right\}\subset\operatorname{co}\left\{f(t)\mid t\in\mathbb{R}\right\}.$
Proof 4.29.
Let $U=\operatorname{co}\{f(t)\mid t\in\mathbb{R}\}$. Define the extended real valued function $I_{U}:\mathbb{R}^{n}\to\mathbb{R}\cup\{-\infty,\infty\}$ as
@@
i.e., $I_{U}(F_{\varepsilon}(t))=0$ so $F_{\varepsilon}(t)\in U$. Since $t$ and $\varepsilon$ are arbitrary, the claim follows.
Remark 4.30.
It is common that the path is a continuous function defined in a compact subset of $\mathbb{R}$, i.e., $f:[a,b]\to\mathbb{R}^{n}$ with $\infty<a<b<\infty$. In such a case, we can extend the function $f$ to $\mathbb{R}$ as follows to get a new continuous function
@@
Note that when working with trajectories with compact domain, we must extend them as done in Remark 4.30.
Lemma 4.32.
Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function and fix $\varepsilon>0$. Let $\bar{f}:[a-\varepsilon,b+\varepsilon]$ be its continuous extension as done in Remark 4.30. The following two statements are true:
@@
Theorem 4.34.
Let $f:[a,b]\to\mathbb{R}^{n}$ be a continuous function and let $\varphi$ be a nonnegative mollifier. Fixed $\varepsilon>0$ let $\bar{f}:[a-\varepsilon,b+\varepsilon]\to\mathbb{R}^{n}$ be the continuous extension as in remark 4.30. Define $F:[a,b]\to\mathbb{R}^{n}$ as $F=\bar{f}*\varphi_{\varepsilon}$; then $L(F)\leq L(f)$.
Proof 4.35.
@@
thus, by definition of the supremum, $L(F)\leq L(f)$.
Remark 4.36.
Note that this does not imply that if $f$ is a geodesic between $f(a)$ and $f(b)$, then $F$ is also a geodesic between these points. This is because $F(a)\neq f(a)$ or $F(b)\neq f(b)$ may occur. While Theorem 4.28 and Remark 4.30 guarantee that $F([a,b])=(\bar{f}*\varphi_{\varepsilon})([a,b])\subset\operatorname{co}f([a,b])$, we cannot ensure that $F$ has the same starting and ending points as $f$. What Theorem 4.34 establishes is that by considering the actual starting and ending points of $F$, we can ensure that $L(F)\leq L(f)$. An example of this property is shown in Figure 5. The original function, which is a linear interpolation of three points in $\mathbb{R}^{2}$ and whose domain is $[a,b]=[0,2]$, is extended to the domain $\varepsilon,2+\varepsilon]$ with $\varepsilon=0.5$ using the extension presented in remark 4.30. As it can be seen from each of its components, $F_{\varepsilon}(a)\neq f(a)$ and $F_{\varepsilon}(b)\neq f(b)$, and clearly $L(F_{\varepsilon})\leq L(f)$ as Theorem 4.34 states. Finally note that, while $f_{2}$ can be considered a geodesic between the points $(0,0)$ and $(2,2)$ in $\mathbb{R}^{2}$ using the Euclidean norm, $(f_{2}*\varphi_{\varepsilon})$ is not a geodesic between those two points.
Figure 5: Visual representation of remark 4.36. The left picture represents the original path $f$ as a blue solid line, and as a red solid line its mollification $F_{\varepsilon}=(f_{1}*\varphi_{\varepsilon},f_{2}*\varphi_{\varepsilon})$ where $\varphi$ is as in (1) and $\varepsilon=0.5$. The middle picture represents the same information but for the first component of the function and its mollification, while the right picture represents the same information but for the second component.
4.4 The effect of reparametrization and mollification
Suppose the desired path is encoded using a continuous function $f:[a,b]\to\mathbb{R}^{n}$ and for a given $\varepsilon>0$ we consider its mollification with parameter $\varepsilon$. As we have seen, we first need to extend the function $f$ to $\bar{f}:[a-\varepsilon,b+\varepsilon]$ as is done in Remark 4.30 or in Theorem 4.34. It may be of interest to reparametrize the curve so it is normalized, i.e., to find a function $g:[-\varepsilon,1+\varepsilon]\to[a-\varepsilon,b+\varepsilon]$, and consider $\bar{f}\circ g*\varphi_{\varepsilon}:[0,1]\to\mathbb{R}^{n}$, as the desired mollified path. How does the parameter of the mollification change under these conditions? That is, does there exist a $\eta=\eta(\varepsilon)>0$ such that
@@ 4.4 The effect of reparametrization and mollification
5 Curvature guarantees from a sequence of waypoints
@@
5.1 The case of three points forming two segments
@@
Definition 5.37 (Two line segments function).
@@
We know by Proposition 4.10 that affine functions are invariant under mollification when Assumption 1 is met and the mollifier is an even function, such as in Example 1. From Figure 3 and Proposition 4.13 it is clear that for the curve defined in Definition 5.37 it happens that $f(1)\neq F_{\varepsilon}(1)$ for all $\varepsilon>0$. However, it of special interesting to compute the intervals in which the original and the mollified curve coincide, because in those intervals there is no approximation, there is an equality between the original and mollified curve. This is easily answered in the next proposition, in which we show that the initial and ending points of the original and mollified curve also coincide.
Proposition 5.38.
Let $\bar{f}$ be as in (6). Suppose $\varepsilon\in(0,\frac{1}{2})$, define for $r\in\{1,2\}$, $V_{r}:=[r-1+\varepsilon,r-\varepsilon]$ and let $F_{\varepsilon}:=\bar{f}*\varphi_{\varepsilon}$ with $\varphi$ even. Then $F_{\varepsilon}|_{V}=\bar{f}|_{V}=f_{V}$, $F_{\varepsilon}(0)=f(0)$, and $F_{\varepsilon}(2)=f(2)$.
Proof 5.39.
Let $t\in V_{r}$, and note that for any $s\in[-\varepsilon,\varepsilon]$ it holds that $r-1\leq t-\varepsilon<t-s<t+\varepsilon\leq r$. Thus, if $r=1$ the function to be mollified is the line $P_{0}+(P_{1}-P_{0})t$ for any $t\in V_{r}$ and if $r=2$ the function to be mollified is the line $P_{1}+(P_{2}-P_{1})(t-1)$ for any $t\in V_{r}$. Thus, by Proposition 4.10 we get the desired result. The proof that the initial and ending points are the same it is carried in a similar fashion.
@@
For $r\in\mathbb{N}$ we define $\tilde{P}_{r}:=P_{r}-P_{r-1}$. Note that both $f$ and $\bar{f}$ are continuous functions, hence locally integrable, and from Theorem 1 we have that if $F_{\varepsilon}:=\bar{f}*\varphi_{\varepsilon}$ where $\varphi$ is any mollifier and $\varepsilon>0$ then $F_{\varepsilon}^{\prime}=\bar{f}*\varphi_{\varepsilon}^{\prime}$. It is not difficult to prove that in this case, $F_{\varepsilon}^{\prime}=\bar{f}*\varphi_{\varepsilon}^{\prime}=\bar{f}^{\prime}*\varphi_{\varepsilon}$ for any $\varepsilon>0$.
We can now exploit these results to obtain a formula for the curvature, as well as an upper bound.
5.1.1 Computing the exact curvature
We know by the discussion above that $F_{\varepsilon}^{\prime}=(\bar{f}*\varphi^{\prime}_{\varepsilon})=(\bar{f}^{\prime}*\varphi_{\varepsilon})$ everywhere. Consider a mollifier $\varphi$ and let $\varepsilon>0$. Given $t\in\mathbb{R}\setminus\{1\}$, note that
@@ 5.1.1 Computing the exact curvature
5.1.2 Upper bounding the curvature
@@
Figure 6: Left plot represents of a three-point-two-segment function and its mollification. Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of (8). In this case the zone of maximum curvature corresponds to a point really close (or equal) to $t=1$. We have used the mollifier presented in Example 1.
Figure 7: Left plot represents of a three-point-two-segment function and its mollification. Right plot represents the curvature of the function and its upper bound $U_{\kappa}$ which is the right hand side of (8). In this case the zone of maximum curvature does not correspond to the join of the two segments. We have used the mollifier presented in Example 1.
5.2 The general case